# APCodec: A Neural Audio Codec with Parallel Amplitude and Phase Spectrum Encoding and Decoding

### v1

@@ III-A3 Decoder
Figure 3: Details of the training losses of the proposed APCodec. Here, VQ, Conv2D and LReLU represent the vector quantizer, 2D convolutional layer and leaky rectified linear unit, respectively. Other abbreviations are the same as those used in Fig. 1. The structure of the encoder, quantizer, and decoder is simplified.
III-B Training Criteria
@@
III-B1 Spectral-level Loss
@@
where $\bm{S}\in\mathbb{C}^{F\times N}$ is the natural short-time complex spectrum extracted from $\bm{x}$. $Re$ and $Im$ are the real part calculation and imaginary part calculation, respectively. It reflects the differences between the decoded short-time complex spectrum and the natural one. To mitigate the inconsistency issues of STFT and narrow the consistency gap between $\hat{\bm{S}}$ and the consistent short-time complex spectrum $\tilde{\bm{S}}=STFT(ISTFT(\hat{\bm{x}}))$, we define the consistency loss as follows:
@@
Ultimately, we articulate the loss on the mel spectrogram as a fusion of MAE and MSE between the extracted mel spectrogram $\hat{\bm{M}}\in\mathbb{R}^{F\times M}$ and $\bm{M}\in\mathbb{R}^{F\times M}$ derived from decoded waveform $\hat{\bm{x}}$ and natural one $\bm{x}$, respectively, i.e.,
$\mathcal{L}_{M}=\dfrac{1}{FM}\cdot\mathbb{E}_{\left(\hat{\bm{M}},\bm{M}\right)}\left(\left\lVert\hat{\bm{M}}-\bm{M}\right\rVert_{1}+\left\lVert\hat{\bm{M}}-\bm{M}\right\rVert_{F}^{2}\right),$ (16)
where $M$ is the dimensionality of the mel spectrogram.
@@ III-B2 Quantization Loss
where $\hat{\bm{C}}_{q}\in\mathbb{R}^{F_{c}\times N_{c}}$ $(1\leq q\leq Q)$ is the output of the $q$-th VQ and $\hat{\bm{C}}_{0}=\bm{C}$. As shown in Fig. 3, the quantization loss $\mathcal{L}_{Q}$ is used exclusively to update the encoder through gradient backpropagation while bypassing the quantizer.
III-B3 GAN-based Loss
@@ III-B4 Training Process
III-C Low-latency Implementation by Knowledge Distillation
@@
To attain low-latency streamable inference, we implement modifications to specific components of the APCodec, covering the following three aspects. 1) Unlike some well-known codecs like SoundStream [25] and Encodec [26] that employ causal convolutions, the streamable APCodec replaces all non-causal convolutional layers (excluding upsampling/downsampling layers) with feed-forward layers. This approach is conducive to reducing model size and improving generation efficiency; 2) Replacing the original non-causal upsampling deconvolutional layers with causal ones; 3) Setting the kernel size of the downsampling convolutional layer to be smaller than the downsampling rate $D$.
@@
However, with the aforementioned modifications, the streamable APCodec, compared to the original non-streamable APCodec, inevitably leads to a deterioration in decoded audio quality. Therefore, we introduce a knowledge distillation training strategy, utilizing a well-trained non-streamable APCodec as the teacher model to guide the training of the streamable APCodec (i.e., the student model). To establish a connection between the teacher model and the student model, we introduce a knowledge distillation loss $\mathcal{L}_{KD}$, defined as the MSE between the features of the two models at corresponding positions. These positions encompass the output of all convolutional layers, feed-forward layers, modified ConvNeXt v2 blocks, and the quantizer in Fig. 1. At the training stage, the streamable APCodec uses $\mathcal{L}+\lambda_{KD}\mathcal{L}_{KD}$ and $\mathcal{L}_{D}$ to train the generator and discriminators alternately, where $\lambda_{KD}$ is hyperparameter. Through training, the streamable APCodec aims to approach the decoded audio quality of the non-streamable APCodec while maintaining its advantage of low latency.
IV Experiments
IV-A Data and Feature Configuration
@@
A subset of the VCTK-0.92 corpus [42] which contained approximately 43 hours of 48 kHz speech recordings from 108 speakers with various accents, was adopted in our experiments. We selected 40,936 utterances from 100 speakers as the training set. Then we built the test set, which included 2,937 utterances from remaining 8 unseen speakers. The original 48 kHz waveforms and downsampled waveforms at 24 kHz and 16 kHz were used for the experiments (i.e., $f_{s}=48000$ or $24000$ or $16000$). When extracting the amplitude spectra, phase spectra and mel spectrograms from natural waveforms, the window size was 320 samples (i.e., $w_{l}=320$), the window shift was 40 samples (i.e., $w_{s}=40$), and the FFT point number was 1024 (i.e., $N=513$). This configuration is applicable to waveforms at all sampling rates.
IV-B Model Details
In our experiments [^1] Source codes will be released after paper acceptance. Examples of generated speech can be found at https://yangai520.github.io/APCodec., we compared the proposed APCodec with four well-known neural waveform codecs, i.e., SoundStream [25], Encodec [26], HiFi-Codec [29] and AudioDec [33]. The descriptions of these codecs are as follows.
@@ IV-B Model Details
  • SoundStream: The SoundStream [25] audio codec. We reimplemented it using the open source implementation [^2] https://github.com/yangdongchao/AcademiCodec.. The downsampling/upsampling ratio was 320. It can achieve low-latency streamable inference.
@@
  • APCodec: The proposed non-streamable APCodec. In the encoder and decoder, the kernel size for all convolutional operations was set to 7. The kernel size for two deconvolutional operations was set to 16. The channel size $C$, $C_{H}$ and $N_{c}$ were 256, 512 and 32, respectively. The downsampling/upsampling ratio was $D=8$. Fig. 1 serves as an example of a 48 kHz audio codec, showcasing the sampling rates of the spectral characteristics and latent codes. We can observe that the APCodec only requires 8$\times$ downsampling to encode a latent code with a sampling rate as low as 150 Hz. The hyperparameters for loss functions were set as $\lambda_{P}=\frac{20}{9}$, $\lambda_{RI}=2.25$, $\lambda_{S}=\frac{4}{9}$, $\lambda_{M}=1$, $\lambda_{spec}=45$, $\lambda_{Q}=7.5$, and $\lambda_{MRD}=0.1$. The model was trained using the AdamW optimizer [43] with $\beta_{1}=0.8$ and $\beta_{2}=0.99$ on a single NVIDIA RTX 3090 GPU. The learning rate decay was scheduled by a 0.999 factor in every epoch with an initial learning rate of 0.0002. The batch size was 16, and the truncated waveform length was 7960 samples for each training step. The model was trained until 1M steps.
  • APCodec-S: The proposed streamable APCodec. It was modified according to the methods outlined in Section III-C for the APCodec, wherein the number of nodes in the replaced feed-forward layers remain consistent with the original convolutional layers’ channel size. The hyperparameter for knowledge distillation was set as $\lambda_{KD}=1$. Other training strategies were consistent with those used in APCodec.
These codecs were comparable because they all employed the similar quantization method (i.e., RVQ or related strategy). All of the above codecs adopted 1024 vectors (i.e., $M=1024$) in the codebook of each VQ. We conducted experiments at all three sampling rates, with two bitrates (low and high) tested at each sampling rate. However, since the AudioDec was originally designed for 48 kHz audio codec, its experiments were conducted only at a sampling rate of 48 kHz. For 48 kHz audio codecs, the bitrates were set at 6 kbps and 12 kbps, respectively. For 24 kHz audio codecs, the bitrates were set at 3 kbps and 6 kbps, respectively. For 16 kHz audio codecs, the bitrates were set at 2 kbps and 4 kbps, respectively. The configuration of low bitrate and high bitrate for SoundStream, Encodec, AudioDec, APCodec and APCodec-S was achieved by setting the number of VQs within the quantizer to $Q=4$ and $Q=8$, respectively. Due to the adoption of the GRVQ quantization strategy in HiFi-Codec, it employed two groups of RVQ, each consisting of 2 and 4 VQs, for achieving encoding for both low and high bitrates, respectively.
IV-C Evaluation Metrics
We comprehensively evaluated the performance of these compared audio codecs using multiple metrics. These metrics were specifically designed to evaluate the amplitude spectrum quality, overall audio quality, intelligibility, phase spectrum quality, generation speed and model complexity, respectively.
@@ IV-C Evaluation Metrics
  • Overall audio quality: The commonly used virtual speech quality objective listener (ViSQOL) [^4] https://github.com/google/visqol. [45] tool was used to assess the overall quality of the decoded audio $\hat{\bm{x}}$, with natural audio $\bm{x}$ as the reference. The ViSQOL outputs a mean opinion score - listening quality objective (MOS-LQO) score, where a higher score indicates better audio quality. The ViSQOL supports only two sampling rates: 48 kHz and 16 kHz. For 48 kHz ViSQOL, the MOS-LQO ranges from 1 to 4.75. For 16 kHz ViSQOL, the MOS-LQO ranges from 1 to 5.
  • Phase spectrum quality: One of the highlights of the proposed APCodec lies in the phase modeling. To validate the effectiveness of phase modeling, the anti-wrapping phase distance (AWPD) proposed in our previous work [46], was employed to evaluate the phase spectrum quality between $\hat{\bm{x}}$ and $\bm{x}$. Similar to the mentioned phase loss in Section III-B1, the AWPD was also computed separately for instantaneous phase, group delay, and instantaneous angular frequency (denoted as $\text{AWPD}_{\text{IP}}$, $\text{AWPD}_{\text{GD}}$ and $\text{AWPD}_{\text{IAF}}$, respectively). After the activation of phase errors using the anti-wrapping function $f_{AW}$, the APDW is calculated in a manner akin to LSD, allowing it to accurately depict the actual phase distortion. A smaller result indicates less distortion.
@@
TABLE I: Experimental results of amplitude spectrum quality, intelligibility, overall audio quality and phase spectrum quality for compared codecs on the test set of the VCTK dataset at three sampling rates and two bitrates. The bold and underline numbers indicate the optimal and sub-optimal results, respectively.
@@
SoundStream Yes 48 kHz 6 kbps 0.937 2.80 0.794 3.11 1.80 1.43 1.45
Encodec Yes 1.04 2.61 0.793 3.31 1.80 1.43 1.47
HiFi-Codec No 0.913 2.44 0.852 3.33 1.80 1.42 1.45
AudioDec Yes 0.847 2.85 0.804 3.98 1.81 1.44 1.46
APCodec No 0.818 1.60 0.875 4.07 1.68 1.40 1.44
@@
SoundStream Yes 48 kHz 12 kbps 0.917 2.50 0.831 3.32 1.79 1.43 1.45
Encodec Yes 0.885 2.17 0.860 3.51 1.79 1.42 1.45
HiFi-Codec No 0.829 1.49 0.941 3.88 1.79 1.40 1.44
AudioDec Yes 0.831 2.31 0.825 4.14 1.81 1.44 1.46
APCodec No 0.796 1.33 0.901 4.26 1.60 1.38 1.42
@@
SoundStream Yes 24 kHz 3 kbps 0.977 3.03 0.804 1.79 1.40 1.44
Encodec Yes 0.958 2.74 0.817 1.79 1.39 1.44
HiFi-Codec No 0.849 2.10 0.875 1.79 1.36 1.44
APCodec No 0.839 2.31 0.856 1.66 1.36 1.42
SoundStream Yes 24 kHz 6 kbps 0.944 2.7 0.832 1.78 1.38 1.44
Encodec Yes 0.933 2.53 0.836 1.78 1.38 1.44
HiFi-Codec No 0.850 1.83 0.910 1.77 1.35 1.43
APCodec No 0.815 2.02 0.877 1.60 1.34 1.41
SoundStream Yes 16 kHz 2 kbps 0.965 3.11 0.804 3.62 1.78 1.36 1.44
Encodec Yes 0.939 2.98 0.810 3.70 1.78 1.36 1.43
HiFi-Codec No 0.910 2.49 0.832 3.84 1.79 1.35 1.43
APCodec No 0.834 2.48 0.852 4.09 1.68 1.33 1.41
SoundStream Yes 16 kHz 4 kbps 0.938 2.76 0.837 3.83 1.76 1.35 1.43
Encodec Yes 0.928 2.78 0.823 3.77 1.77 1.35 1.43
HiFi-Codec No 0.875 2.14 0.869 4.10 1.77 1.33 1.42
APCodec No 0.792 2.12 0.885 4.32 1.56 1.29 1.38
IV-D Primary Experimental Results

### v2

@@ II-A3 Decoder
Figure 3: Details of the training losses of the proposed APCodec. Here, VQ, Conv2D and LReLU represent the vector quantizer, 2D convolutional layer and leaky rectified linear unit, respectively. MSE, MAE, AW-IP, AW-GD and AW-IAF represent mean square error, mean absolute error, anti-wrapping instantaneous phase, anti-wrapping group delay and anti-wrapping instantaneous angular frequency, respectively. STFT and ISTFT represent the short-time Fourier transform and inverse short-time Fourier transform, respectively. The structure of the encoder, quantizer, and decoder is simplified.
II-B Training Criteria
@@
II-B1 Spectral-level Loss
@@
where $\bm{S}\in\mathbb{C}^{F\times N}$ is the natural short-time complex spectrum extracted from $\bm{x}$. $Re$ and $Im$ are the real part calculation and imaginary part calculation, respectively. It reflects the differences between the decoded short-time complex spectrum and the natural one. To mitigate the inconsistency issues of STFT and narrow the consistency gap between $\hat{\bm{S}}$ and the consistent short-time complex spectrum $\tilde{\bm{S}}=STFT(ISTFT(\hat{\bm{S}}))$, we define the consistency loss as follows:
@@
Ultimately, we articulate the loss on the mel spectrogram as a fusion of MAE and MSE between the extracted mel spectrogram $\hat{\bm{M}}\in\mathbb{R}^{F\times N_{mel}}$ and $\bm{M}\in\mathbb{R}^{F\times N_{mel}}$ derived from $\hat{\bm{x}}$ and $\bm{x}$, respectively, i.e.,
$\mathcal{L}_{M}=\dfrac{1}{FN_{mel}}\cdot\mathbb{E}_{\left(\hat{\bm{M}},\bm{M}\right)}\left(\left\lVert\hat{\bm{M}}-\bm{M}\right\rVert_{1}+\left\lVert\hat{\bm{M}}-\bm{M}\right\rVert_{F}^{2}\right),$ (16)
where $N_{mel}$ is the dimensionality of the mel spectrogram.
@@ II-B2 Quantization Loss
The quantization loss $\mathcal{L}_{Q}$ updates the parameters of the encoder and quantizer separately through gradient detachment operation.
II-B3 GAN-based Loss
@@ II-B4 Training Process
II-C Low-latency Implementation by Knowledge Distillation
@@
To attain low-latency streamable inference, we implement modifications to specific components of the APCodec, covering the following three aspects. 1) Unlike some well-known codecs like SoundStream [25] and Encodec [26] that employ causal convolutions, the streamable APCodec replaces all non-causal convolutional layers (excluding upsampling/downsampling layers) with feed-forward layers. This approach is conducive to reducing model size and improving generation efficiency; 2) Replacing the original non-causal upsampling deconvolutional layers with causal ones; 3) Setting the kernel size of the downsampling convolutional layer to be smaller than or equal to $12D-$ 1. For the last aspect, we provide a detailed explanation as follows. To ensure the generation of at least one frame of latent code, the minimum length of the input audio for APCodec is $w_{s}\cdot D$ samples (i.e., fixed latency). Therefore, the input features of the downsampling convolutional layer have at least $D$ frames. During the convolution operation, $D-1$ zeros are padded before the features. Hence, downsampling convolutional operations can be performed without utilizing future information when the kernel size is smaller than or equal to $12D-$ 1.
@@
However, with the aforementioned modifications, the streamable APCodec, compared to the original non-streamable APCodec, inevitably leads to a deterioration in decoded audio quality. Therefore, we introduce a knowledge distillation training strategy, utilizing a well-trained non-streamable APCodec as the teacher model to guide the training of the streamable APCodec (i.e., the student model). To establish a connection between the teacher model and the student model, we introduce a knowledge distillation loss $\mathcal{L}_{KD}$, defined as the MSE between the features of the two models at corresponding positions. These positions encompass the output of all convolutional layers, feed-forward layers, modified ConvNeXt v2 blocks, and the quantizer in Fig. 1. At the training stage, the streamable APCodec uses $\mathcal{L}+\lambda_{KD}\mathcal{L}_{KD}$ and $\mathcal{L}_{D}$ to train the generator and discriminators alternately, where $\lambda_{KD}$ is hyperparameter. The other hyperparameters used for training the streamable APCodec, as well as the dataset and the total training steps, are entirely consistent with training the non-streamable APCodec. Through training, the streamable APCodec aims to approach the decoded audio quality of the non-streamable APCodec while maintaining its advantage of low latency.
III Experiments
III-A Data and Feature Configuration
@@
A subset of the VCTK-0.92 corpus [43] which contained approximately 43 hours of 48 kHz speech recordings from 108 speakers with various accents, was adopted in our experiments. We selected 40,936 utterances from 100 speakers as the training set. Then we built the test set, which included 2,937 utterances from remaining 8 unseen speakers. The original 48 kHz waveforms and downsampled waveforms at 24 kHz and 16 kHz were used for the experiments (i.e., $f_{s}=48000$ or $24000$ or $16000$). When extracting the amplitude spectra, phase spectra and mel spectrograms from natural waveforms, the window size was 320 samples (i.e., $w_{l}=320$), the window shift was 40 samples (i.e., $w_{s}=40$), and the FFT point number was 1024 (i.e., $N=513$). The dimensionality of the mel spectrograms was 80 (i.e., $N_{mel}=80$). This configuration is applicable to waveforms at all sampling rates.
III-B Model Details
In our experiments [^1] Source codes are available at https://github.com/yangai520/APCodec. Examples of generated audio can be found at https://yangai520.github.io/APCodec., we constructed non-streamable and streamable APCodec models to fairly compare with existing non-streamable and streamable codec models, respectively. The descriptions of the non-streamable and streamable APCodec are as follows.
  • APCodec: The proposed non-streamable APCodec. In the encoder and decoder, the kernel size for all convolutional operations was set to 7. The kernel size for two deconvolutional operations was set to 16. The channel size $K$, $K_{H}$ and $N_{c}$ were 256, 512 and 32, respectively. The downsampling/upsampling ratio was $D=8$. Fig. 1 serves as an example of a 48 kHz audio codec, showcasing the frame rates of the spectral characteristics and latent codes. We can observe that the APCodec only requires 8$\times$ downsampling to encode a latent code with a frame rate as low as 150 Hz. The hyperparameters for loss functions were set as $\lambda_{P}=\frac{20}{9}$, $\lambda_{RI}=2.25$, $\lambda_{S}=\frac{4}{9}$, $\lambda_{M}=1$, $\lambda_{spec}=45$, $\lambda_{Q}=7.5$, and $\lambda_{MRD}=0.1$. The model was trained using the AdamW optimizer [44] with $\beta_{1}=0.8$ and $\beta_{2}=0.99$ on a single NVIDIA RTX 3090 GPU. The learning rate decay was scheduled by a 0.999 factor in every epoch with an initial learning rate of 0.0002. The batch size was 16, and the truncated waveform length was 7960 samples for each training step. The model was trained until 1M steps.
  • APCodec-S: The proposed streamable APCodec. It was modified according to the methods outlined in Section II-C for the APCodec, wherein the number of nodes in the replaced feed-forward layers remain consistent with the original convolutional layers’ channel size. The kernel size of the downsampling convolutional layers was set to 7. It is trained with guidance from the well-trained APCodec. The hyperparameter for knowledge distillation was set as $\lambda_{KD}=1$. Other training strategies were consistent with those used in APCodec.
For high-sampling-rate audio coding, we compared the proposed APCodec with the following codecs:
@@ III-B Model Details
  • Encodec: The Encodec [26] audio codec. It supports audio coding at sampling rates of 24 kHz and 48 kHz, as reported in [26]. We reimplemented it using the open source implementation [^2] https://github.com/yangdongchao/AcademiCodec.. The downsampling/upsampling ratio was 320. It can achieve low-latency streamable inference.
  • DAC: The DAC [34] audio codec. It is designed for 44.1 kHz audio coding. We reimplemented it using the open source implementation [^4] https://github.com/descriptinc/descript-audio-codec. and applied it to 48 kHz audio coding. The downsampling/upsampling ratio for the model was 320. However, the DAC is non-streamable, and there is no streamable implementation provided in the open-source code.
Although our proposed APCodec was initially designed for 48 kHz audio coding, to ensure fair comparisons with some low-sampling-rate audio codecs, we also conducted experiments at lower sampling rates, such as 16 kHz and 24 kHz. In addition to Encodec, the low-sampling-rate audio codecs used for comparison also included the following:
@@
These codecs were comparable because they all employed the similar quantization method (i.e., RVQ or related strategy). All of the above codecs adopted 1024 vectors (i.e., $M=1024$) in the codebook of each VQ. We conducted experiments at all three sampling rates, with two bitrates (low and high) tested at each sampling rate. For 48 kHz audio codecs, the bitrates were set at 6 kbps and 12 kbps, respectively. For 24 kHz audio codecs, the bitrates were set at 3 kbps and 6 kbps, respectively. For 16 kHz audio codecs, the bitrates were set at 2 kbps and 4 kbps, respectively. The configuration of low bitrate and high bitrate for APCodec, APCodec-S, Encodec, AudioDec, DAC and SoundStream was achieved by setting the number of VQs within the quantizer to $Q=4$ and $Q=8$, respectively. Due to the adoption of the GRVQ quantization strategy in HiFi-Codec, it employed two groups of RVQ, each consisting of 2 and 4 VQs, for achieving audio coding for both low and high bitrates, respectively.
III-C Evaluation Metrics
First, we comprehensively evaluated the performance of these compared audio codecs using multiple objective metrics. These objective metrics were specifically designed to evaluate the amplitude spectrum quality, overall audio objective quality, intelligibility, phase spectrum quality, generation speed and model complexity, respectively.
@@ III-C Evaluation Metrics
  • Overall audio objective quality: The commonly used virtual speech quality objective listener (ViSQOL) [^5] https://github.com/google/visqol. [46] tool was used to objectively assess the overall quality of the decoded audio $\hat{\bm{x}}$, with natural audio $\bm{x}$ as the reference. The ViSQOL outputs a mean opinion score - listening quality objective (MOS-LQO) score, where a higher score indicates better audio quality. The ViSQOL supports only two sampling rates: 48 kHz and 16 kHz. For 48 kHz ViSQOL, the MOS-LQO ranges from 1 to 4.75. For 16 kHz ViSQOL, the MOS-LQO ranges from 1 to 5. It should be noted that for the assessment of audio quality at a 24 kHz sampling rate, we upsampled both the decoded audio and reference audio to 48 kHz, and then calculated MOS-LQO using ViSQOL’s 48 kHz mode.
  • Phase spectrum quality: One of the highlights of the proposed APCodec lies in the phase modeling. To validate the effectiveness of phase modeling, the anti-wrapping phase distance (AWPD) proposed in our previous work [47], was employed to evaluate the phase spectrum quality between $\hat{\bm{x}}$ and $\bm{x}$. Similar to the mentioned phase loss in Section II-B1, the AWPD was also computed separately for instantaneous phase, group delay, and instantaneous angular frequency (denoted as $\text{AWPD}_{\text{IP}}$, $\text{AWPD}_{\text{GD}}$ and $\text{AWPD}_{\text{IAF}}$, respectively). After the activation of phase errors using the anti-wrapping function $f_{AW}$, the AWPD is calculated in a manner akin to LSD, allowing it to accurately depict the actual phase distortion. A smaller result indicates less distortion.
@@
Furthermore, to assess human perception of the decoded audio quality, we also conducted subjective experiments. Since the focus of this paper is on high-sampling-rate audio coding, subjective experiments were conducted only on the 48 kHz sampling rate configuration. We conducted a MUSHRA (MUltiple Stimuli with Hidden Reference and Anchor) test [48] on the crowdsourcing platform Amazon Mechanical Turk [^6] https://www.mturk.com. to evaluate the quality of the 48 kHz audio decoded by APCodec, APCodec-S, Encodec, AudioDec and DAC at 6 kbps and 12 kbps on the test set of the VCTK dataset. 20 test utterances decoded by each experimental model were evaluated by about 40 English native listeners. Listeners were asked to give a score between 0 and 100 to each test sample (the reference natural audio tracks had a maximum score of 100).
TABLE I: Objective experimental results for compared codecs on the test set of the VCTK dataset at three sampling rates and two bitrates. The bold and underline numbers indicate the optimal and sub-optimal results, respectively.
@@
APCodec No 48 kHz 6 kbps 0.818 1.60 0.875 4.07 1.68 1.40 1.44
@@
Encodec [26] Yes 1.04 2.61 0.793 3.31 1.80 1.43 1.47
AudioDec [33] Yes 0.847 2.85 0.804 3.98 1.81 1.44 1.46
DAC [34] No 0.841 1.87 0.906 3.81 1.78 1.40 1.44
APCodec No 48 kHz 12 kbps 0.796 1.33 0.901 4.26 1.60 1.38 1.42
@@
Encodec [26] Yes 0.885 2.17 0.860 3.51 1.79 1.42 1.45
AudioDec [33] Yes 0.831 2.31 0.825 4.14 1.81 1.44 1.46
DAC [34] No 0.815 1.76 0.954 4.06 1.75 1.38 1.42
APCodec No 24 kHz 3 kbps 0.839 2.31 0.856 4.08 1.66 1.36 1.42
APCodec-S Yes 0.864 2.18 0.838 4.11 1.78 1.38 1.44
Encodec [26] Yes 0.958 2.74 0.817 3.82 1.79 1.39 1.44
SoundStream [25] Yes 0.977 3.03 0.804 3.79 1.79 1.40 1.44
HiFi-Codec [29] No 0.849 2.10 0.875 4.05 1.79 1.36 1.44
APCodec No 24 kHz 6 kbps 0.815 2.02 0.877 4.28 1.60 1.34 1.41
APCodec-S Yes 0.812 1.64 0.889 4.35 1.66 1.35 1.42
Encodec [26] Yes 0.933 2.53 0.836 3.81 1.78 1.38 1.44
SoundStream [25] Yes 0.944 2.70 0.832 3.90 1.78 1.38 1.44
HiFi-Codec [29] No 0.850 1.83 0.910 4.13 1.77 1.35 1.43
APCodec No 16 kHz 2 kbps 0.834 2.48 0.852 4.09 1.68 1.33 1.41
APCodec-S Yes 0.856 2.56 0.851 4.05 1.73 1.33 1.42
Encodec [26] Yes 0.939 2.98 0.810 3.70 1.78 1.36 1.43
SoundStream [25] Yes 0.965 3.11 0.804 3.62 1.78 1.36 1.44
HiFi-Codec [29] No 0.910 2.49 0.832 3.84 1.79 1.35 1.43
APCodec No 16 kHz 4 kbps 0.792 2.12 0.885 4.32 1.56 1.29 1.38
APCodec-S Yes 0.810 1.88 0.881 4.35 1.66 1.32 1.41
Encodec [26] Yes 0.928 2.78 0.823 3.77 1.77 1.35 1.43
SoundStream [25] Yes 0.938 2.76 0.837 3.83 1.76 1.35 1.43
HiFi-Codec [29] No 0.875 2.14 0.869 4.10 1.77 1.33 1.42
III-D Primary Experimental Results